OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the circumference of a circle, C = 2πr A: C/r= 2π B: C/π= 2r C: C/rπ= 2 D: Cr=2π

OpenStudy (anonymous):

C i am sure

OpenStudy (unklerhaukus):

\[w=xyz\] \[\frac wx =\frac{xyz}x\] \[\frac wx =yz\]

OpenStudy (ksaimouli):

D is correct

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):

this is d but they gave another one so D is correct answer for ur Q

OpenStudy (anonymous):

D isn't in a fraction form its suppose to be like this Cr.. Honestly I'm not completely sure what any of the answers are. I've been stuck on this question for forever haha

OpenStudy (unklerhaukus):

thats right to get one off the terms off the right hand side we divide both sides of the equation by that term